### v1

# Information Flow Routes: Automatically Interpreting Language Models at Scale

Abstract
Information flows by routes inside the network via mechanisms implemented in the model. These routes can be represented as graphs where nodes correspond to token representations and edges to operations inside the network. We automatically build these graphs in a top-down manner, for each prediction leaving only the most important nodes and edges. In contrast to the existing workflows relying on activation patching, we do this through attribution: this allows us to efficiently uncover existing circuits with just a single forward pass. Additionally, the applicability of our method is far beyond patching: we do not need a human to carefully design prediction templates, and we can extract information flow routes for any prediction (not just the ones among the allowed templates). As a result, we can talk about model behavior in general, for specific types of predictions, or different domains. We experiment with Llama 2 and show that the role of some attention heads is overall important, e.g. previous token heads and subword merging heads. Next, we find similarities in Llama 2 behavior when handling tokens of the same part of speech. Finally, we show that some model components can be specialized on domains such as coding or multilingual texts.
Machine Learning, ICML
${}^{\ast}$ork done during an internship at Meta AI.
1 Introduction
Current state-of-the-art LMs are built on top of the Transformer architecture (Vaswani et al., 2017; Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a, b). Inside the model, each representation evolves from the current input token embedding [^1] Sometimes, along with positional encoding. to the final representation used to predict the next token. This evolution happens through additive updates coming from attention and feed-forward blocks. The resulting stack of same-token representations is usually referred to as “residual stream” (Elhage et al., 2021), and the overall computation inside the model can be viewed as a sequence of residual streams connected through layer blocks. Formally, we can see it as a graph where nodes correspond to token representations and edges correspond to operations inside the model (attention heads, feed-forward layers, etc.).
Figure 1: The important information flow routes for a token (Mary) prediction. GPT2-Small, $\tau=0.04$.
While during model computation all the edges (i.e., model components) are present, computations important for each prediction are likely to form only a small portion of the original graph (Voita et al. (2019); Wang et al. (2023); Hanna et al. (2023), among others). We extract this important subgraph in a top-down manner by tracing information back through the network and, at each step, leaving only edges that were relevant (Figure 1). To understand which edges are important, we rely on an attribution method (Ferrando et al., 2022) and refuse from activation patching [^2] Originally introduced by Vig et al. (2020) to analyze LMs through the lens of causal mediation analysis. typical for the existing mechanistic interpretability workflows (Wang et al., 2023; Hanna et al., 2023; Conmy et al., 2023; Stolfo et al., 2023; Heimersheim & Janiak, 2023). Firstly, patching requires human efforts to create templates and contrastive examples and, therefore, is only applicable to a few pre-defined templates. Secondly, explaining a single prediction demands a substantial number of interventions (patches). Given that each intervention needs a forward pass, studying large models becomes increasingly impractical. In contrast, our method is about 100 times faster.
In the experiments, we first show that our information flow routes rely on the same task-specific attention heads found in patching circuits (Wang et al., 2023; Hanna et al., 2023). However, our method is more versatile and informative than patching: it can evaluate the importance of model components both (i) overall for a prediction, and (ii) compared to a contrastive example (i.e., patching setting). What is more, we argue that patching is fragile: its results can vary depending on the choice of the contrastive template (i.e., human preference).
Figure 2: Full information flow graph.
Next, we come to the settings unreachable to patching, i.e. the broad set of predictions and general importance. We experiment with Llama 2 and show that some attention head functions are overall important, e.g. previous token heads and subword merging heads. Next, we find that information inside Llama 2 flows similarly when handling tokens of the same part of speech. Finally, we show that some model components are specialized on domains such as coding or multilingual texts: they are active for these domains and not active otherwise.
Overall, our contributions are as follows:
  • we propose to explain predictions of transformer LMs via information flow routes;
  • compared to patching circuits, our method is (i) applicable to any prediction, (ii) more informative, and (iii) around 100 times faster;
  • we analyze the information flow of Llama 2 and find model components that are (i) generally important, and (ii) specific to domains.
2 Extracting Information Flow Routes
Figure 2 illustrates computations inside a Transformer LM along with the intermediate representations after each block.
Figure 3: General-case algorithm for extracting the important subgraph of an information flow graph.
Information flow graph.
In this view, we can see a graph where (i) nodes correspond to token representations and (ii) edges correspond to operations inside the network that move information across nodes. Specifically, $x_{pos}^{lA}$ and $x_{pos}^{l}$ are representations of the token at position $pos$ after applying only attention block in layer $l$ or the entire layer, respectively. Each of $x_{1}^{l-1}$, …, $x_{pos}^{l-1}$ is connected to $x_{pos}^{lA}$ via attention edges (and via a residual stream edge from $x_{pos}^{l-1}$ to $x_{pos}^{lA}$), $x_{pos}^{lA}$ is connected to $x_{pos}^{l}$ via two edges: via feed-forward layers (FFNs) and via residual stream.
Extracting the important subgraph.
While during model computation all the edges from Figure 2 are present, computations important for each prediction are likely to form only a small portion of the original graph (Voita et al. (2019); Wang et al. (2023); Hanna et al. (2023), etc). We extract this important subgraph in a top-down manner by tracing information back through the network (Figure 3). We start building the important subgraph from a single node – the representation on top of the residual stream. Then, we go over immediately connected lower nodes and, if they were important at this step (i.e., contributed sufficiently), add them to the graph along with the corresponding edges. Note that for $x_{pos}^{l}$, immediate connections contain $x_{pos}^{lA}$ with two different edges: via FFN and via residual connection, and we treat these connections separately.
Importance via attribution, not patching.
In the algorithm above, what is left is to define how to compute the importance of an edge. While lately it has become rather typical to use patching (Wang et al., 2023; Hanna et al., 2023; Conmy et al., 2023; Stolfo et al., 2023; Heimersheim & Janiak, 2023), instead of patching, we choose to use attribution. This choice is crucial for our work and makes our method about 100 times faster than alternatives while (i) being able to recover previously discovered circuits, (ii) doing this in a more versatile manner, and (iii) leading to many new observations.
In the rest of this section, we explain the specific attribution method we use.
2.1 Evaluating Edge Importance
For the attribution method, we adopt ideas from ALTI (Aggregation of Layer-Wise Token-to-Token Interactions) by Ferrando et al. (2022). While originally this method propagated attributions from output to input tokens through the entire model, we take only its part that defines contributions between immediately connected nodes. We choose this method because it is straightforward, easy to implement, and was shown effective for practical applications, e.g. detecting hallucinations in neural machine translation (Dale et al., 2023a, b; Guerreiro et al., 2023).
2.1.1 Importance in General Case
Note that in our graph (Figure 2), each node can be represented as a sum of the vectors corresponding to the incoming edges. According to ALTI, importance of one of the vectors (edges) to the overall sum (node) is proportional to its proximity to the resulting sum. Formally, if
${\bm{x}}={\bm{y}}_{1}+\dots+{\bm{y}}_{m},\vspace{-2ex}$
then
$\!\!\!\!importance({\bm{y}}_{j},{\bm{x}})\!=\!\frac{proximity({\bm{y}}_{j},{\bm{x}})}{\sum\limits_{k}proximity({\bm{y}}_{k},{\bm{x}})},$ (1)
$proximity({\bm{y}}_{j},{\bm{x}})=\max(-||{\bm{y}}_{j}-{\bm{x}}||_{1}+||{\bm{x}}||_{1},0).$
Here, we use negative distance as a measure of similarity: the smaller the distance between ${\bm{y}}_{i}$ and ${\bm{x}}$, the higher the impact of ${\bm{y}}_{i}$ on ${\bm{x}}$. Note also that we ignore contributions of the vectors lying beyond the $l_{1}$ length of ${\bm{x}}$. For more details, see Ferrando et al. (2022).
Now, we need to define vector updates corresponding to the edges for the FFN and the attention blocks.
2.1.2 Defining FFN Edges
For the FFN blocks, edge vectors are straightforward. In the notation of Figure 2,
${\bm{x}}_{pos}^{l}$ $={\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\bm{x}}_{pos}^{lA}}+{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\text{FFN}_{l}({\bm{x}}_{pos}^{lA})},$
and the two terms correspond to the edges via residual connection and via FFN, respectively.
Figure 4: Decomposition of an update coming from an attention head into per-input terms. Layer indices are omitted for readability.
2.1.3 Defining Attention Edges
Attention edges correspond to connections between residual streams. To get vector updates coming from other residual streams to the current node, we follow previous work and decompose the output of an attention block into sum of vectors corresponding to its inputs (Kobayashi et al., 2020; Ferrando et al., 2022). We illustrate this decomposition in Figure 4.
Individual attention heads.
Formally, for attention head $h$
$Attn^{h}(\mathbf{x}_{\leq i})=\sum_{j\leq i}\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j}),$ (2)
where $f^{h}({\bm{x}}_{j})={\bm{x}}_{j}LW^{h}_{OV}$, $W_{OV}^{h}=W_{V}^{h}W_{O}^{h}$ the values and output combined matrix for head $h$, $\alpha_{i,j}^{h}$ are scalar attention weights, and $L$ is the linearized layer normalization (see Appendix A). [^3] From here onwards, we omit layer indices for readability.
In a bit more detail, a typical attention implementation (1) weights corresponding value vectors within each attention head, (2) concatenates head outputs, and (3) further multiplies by the output matrix $W_{O}$. In our view above, we split the output matrix $W_{O}$ into its head-specific parts $W_{O}^{h}$, drag these parts inside attention heads, and combine them with the head’s values matrix: $W_{OV}^{h}=W_{V}^{h}W_{O}^{h}$.
Overall, Figure 4 shows that on the way from attention head input to its output, each representation ${\bm{x}}_{j}$ is transformed linearly into ${f^{h}({\bm{x}}_{j})={\bm{x}}_{j}LW_{OV}^{h}}$ and multiplied by a scalar attention weight $\alpha_{i,j}^{h}$. This gives us the “raw output” emitted by each input vector ${\bm{x}}_{j}$ when treating attention weights as prediction-specific constants. In this view, information flows through attention heads by independent channels $\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j})$ that converge in the next residual stream state.
Figure 5: Decomposition of an update coming from an entire attention layer into per-input terms. Layer indices are omitted for readability.
Entire attention block.
Figure 5 shows information flow for the entire attention block. Similarly to individual attention heads, here information also flows via independent channels corresponding to input tokens:
$Attn(\mathbf{x}_{\leq i})=\sum\limits_{h}^{H}\sum_{j\leq i}\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j}).$ (3)
Following the general view from Section 2.1.1, we compute the importance of each of the terms in this sum individually and then aggregate edge capacities connecting the same pair of nodes. That is, sum over heads within each token, and additionally include a residual connection for the current token (Figure 5). Formally, the final attention edge importances are as follows:
${e_{i,j}=\left\{\begin{array}[]{@{}ll@{}}\sum_{h}^{H}e^{h}_{i,j}&\mbox{if}~{}j\neq i\\ \sum_{h}^{H}e^{h}_{i,j}+e^{r}_{i}&\mbox{if}~{}j=i\end{array}\right.}$
where $e_{i,j}^{h}$ are token-head importances and $e_{i}^{r}$ is the residual connection importance. Formally, ${e^{h}_{i,j}=importance(\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j}),{\bm{x}}^{A}_{i})}$ and ${e^{r}_{i}=importance({\bm{x}}_{i},{\bm{x}}^{A}_{i})}$ respectively, as defined in equation (1).
2.2 Extracting the Important Subgraph
In the previous Section 2.1 we explained how we compute the importances of the edges in the full information flow graph (Figure 2) that are later used in Algorithm 3 to extract the important subgraph. The algorithm requires setting a threshold $\tau$ for the minimum edge importance in the graph which will complete the algorithm. One can make this algorithm more fine-grained by taking into account a few considerations.
Minimal head importance requirement.
First, we can exclude attention edges composed only of unimportant heads. Intuitively, if between two nodes none of the attention heads contributed significantly, then this connection might be considered unimportant. In this case, in equation (4) we set to zero the connections with edge values below $\tau$ and then renormalize the rest. Either of the two versions of the algorithm (with zeroing-out unimportant sub-edges or not) can be used – in our experiments, we did not notice significant differences. In the main text, we show results with this strategy.
Different thresholds for edge types.
Generally, feed-forward blocks have higher importance than attention heads. This is expected: while attention heads are only a (small) part of the attention block, the feed-forward block is not decomposed further. Therefore, one can set different thresholds for retaining attention and FFN edges. However, we do not do this in our experiments.
Figure 6: Greater-than, GPT2-Small. Attention head activation frequency ($\tau=0.03$).
Figure 7: IOI, GPT2-Small. Attention head activation frequency ($\tau=0.03$).
3 Information Flow vs Patching Circuits
First, we compare the information routes extracted with our method to the circuits found in previous work for two tasks: Indirect Object Identification (IOI) (Wang et al., 2023) and Greater-than tasks (Hanna et al., 2023). The IOI task consists of predicting the next word in sentences like “When Mary and John went to the store, John gave a drink to ___”. An initial clause features the names of two individuals (Mary and John). Then, the second clause depicts a person exchanging an item with the other person. The goal is to predict the second name in the main clause, i.e. the Indirect Object (Mary). In the Greater-than task, the model is prompted to predict a number given a sentence following the template: “The [noun] lasted from the year XXYY to the year XX___”. Here, the task is to predict a number higher than YY.
Information flow vs patching.
The circuits for these tasks were previously found using activation patching and relying on “baseline” (contrastive) templates. These templates are designed to bring to the surface the specific task. For example, for IOI a contrastive template contains three different names instead of repeating one of them (Figure 7). One of the differences between information flow and patching results is that our method finds all the components that contributed to the prediction, while patching is claimed to find what is important for the original task but not the contrastive baseline. As we will see, (i) our method can also be used in this manner, and (ii) it gives more reasonable results.
3.1 Indirect Object Identification
For the IOI task, the previously found circuit contains several attention heads with various functions, e.g. Name Mover Heads, Duplicate Token Heads, etc. (Wang et al., 2023).
Using our method, we extract information flow routes for both the original task and the contrastive templates, and evaluate the activation frequency of attention heads. When looking at overall contributions, Figure 7 (left) shows that our routes largely consist of the heads discovered previously. There are, however, several heads which are always part of the routes (have near 1 activation frequency) but are not in the “patching circuit” (shown with pale circles). Interestingly, when looking at the difference with contrastive templates (Figure 7, right), these generic heads disappear. This makes sense: information flow routes contain all the components that were important for prediction, and these contain (i) generic components that are overall important, and (ii) task-specific heads. Note also how our method goes further than patching for identifying the difference between the original and the contrastive tasks. Figure 7 shows that e.g. the previous token heads (yellow) are important for both types of predictions and are not specific to the IOI task.
3.2 Greater-than
For the overall contributions in the greater-than task (Figure 6, left), we also see that (i) task-specific heads discovered via patching are among the most important heads in the information flow routes, and (ii) many other heads are important. As before, let us now look at the differences with the contrastive templates.
Contrastive template matters.
Interestingly, when using the same template as Hanna et al. (2023), i.e. “…from the year XXYY” with $\text{YY}=01$, instead of $\text{YY}>01$ overall contributions do not change much: all heads important for the original task are also important for the contrastive template (Figure 6, center). While this contradicts the original work by Hanna et al. (2023), this is expected: predicting a number higher than “01” still requires greater-than reasoning (not all numbers would fit after “01”, e.g. “00” would not). In contrast, if we consider another template with $\text{YY}=00$, overall contributions change and “greater-than heads” (red in Figure 6) become less important. This difference in the results for “01” vs “00” in the contrastive template highlights the fragility of patching: not only it requires human-defined contrastive templates, but also patching results are subjective since they vary depending on the chosen template.
Overall, we see that, compared to patching, information flow routes are more versatile and informative. Indeed, they can find the importance of model components both (i) overall for a prediction, and (ii) compared to a contrastive example (i.e., patching setting), and are able to show differences between the contrastive templates.
3.3 OPT-125m vs GPT2-Small
Additionally, we experiment with OPT-125m which has the same number of layers and heads than GPT2-small. We find that the information flow routes for IOI and greater-than tasks are similar for OPT-125M and GPT2-small (Appendix B).
3.4 Hundred Times Faster than Patching
Obtaining information flow routes involves a two-step process. First, we run a forward pass and cache internal activations. Then, we obtain edge importances to build the subgraph as depicted in Algorithm 3. For comparison purposes, we contrast our approach with the ACDC algorithm introduced by (Conmy et al., 2023), an automated circuit discovery algorithm based on patching. According to their findings, on the IOI task with a batch of 50 examples, it requires 8 minutes on a single GPU to discover the circuit. In contrast, our method accomplishes this task in 5 seconds, around 100x in time reduction.
Figure 8: t-SNE of component importance vectors. Coloured by: (a) input token POS Tag, (b) next token POS tag, (c) whether the input token is the first or a later subword. Llama 2-7B.
4 General Experiments
Information flows inside the network via mechanisms implemented in the model. In this section, we look at general patterns in component importances, and try to understand if important mechanisms depend on the type of tokens processed.
Setting.
Using the full information flow graph (Figure 2) of Llama 2-7B, we look at all the immediate connections with each residual stream. We then record the importance values of all the sub-edges corresponding to individual attention heads (i.e., $e_{i,j}^{h}$), as well as FFN blocks. We use a subset of 1000 sentences from the C4 dataset (Raffel et al., 2020).
4.1 Component Importance for POS
First, let us see how component importance depends on parts of speech (POS). For this, we pack per-prediction component importances into vectors [^4] Each vector corresponding to the $i$-th position is defined as $\left(\sum_{j}e^{1,1}_{i,j},\sum_{j}e^{1,2}_{i,j},\ldots,\sum_{j}e^{L,H}_{i,j},e^{\mbox{ffn}_{1}}_{i},\ldots,e^{\mbox{ffn}_{L}}_{i}\right)$. and apply t-SNE (van der Maaten & Hinton, 2008). Figure 8 shows the resulting projection with datapoints colored according to either input or next token [^5] Here, we take the next token from the dataset and not the one generated by the model. While this adds some noise to the results, we expect at least parts of speech of the reference and the predicted tokens to be similar. part of speech. [^6] We obtain part of speech tags using NLTK library (universal tagset, Petrov et al. (2012)) at the word level and assign a tag to all subwords of the word.
Content vs function words.
From Figure 8a we see that for function words as inputs, component contributions are clustered according to their part of speech tag. Roughly speaking, for these parts of speech the model has “typical” information flow routes and, by knowing component contributions, we can infer input token POS rather accurately. Interestingly, this is not the case for content words: while we can see verbs being somewhat separated from nouns, overall contribution patterns for content words are mixed together in a large cluster. Apparently, the reasoning for these words is more complicated and is defined by a broader context rather than simply input token POS.
First vs later subwords.
Looking in a bit more detail at the input tokens, Figure 8c shows that contribution patterns strongly depend on whether the current token is the first or a later subword of some word. Clearly, there are some model components specialized for first or later subwords – we confirm this later in Section 4.3. Note also that Figure 8c explains the two distinct clusters for numbers we saw in Figure 8a (purple): they separate number tokens into first and later digits.
Patterns wrt to current vs next token.
Finally, Figure 8b shows the same datapoint colored by part of speech of the next token. Comparing this to Figure 8a, we see that contribution patterns depend more on input tokens rather than output tokens. This might be because in the lower part of the network, a model processes inputs in a generic manner, while the higher network part (that prepares to generate output) is more fine-grained and depends on more attributes than part of speech.
[…]
4.2 Bottom-to-Top Patterns
Since we already started talking about functions of the lower and the higher parts of the network, let us now look at the bottom-up contribution patterns. Figure 9a shows as average activation frequency of attention heads for a small $\tau=0.01$ – this gives us an estimate “of” how many times an attention head affects a residual stream. Additionally, we show the importance of the FFN block in each layer. As expected, attention and feed-forward blocks are more active at the bottom part of the network, where the model processes input and extracts general information. Later, when a model performs more fine-grained and specialized operations, the activation frequency of individual attention heads and FFN blocks goes down. Interestingly, the last-layer FFN is highly important: apparently, this last operation between the residual stream and the unembedding matrix (i.e., prediction) changes representation rather significantly.
4.3 Positional and Subword Merging Heads
Some of our observations above hint at certain attention head functions that might be overall important. First, in Section 3.1 we showed that previous tokens heads are generally important in the IOI task, both for target examples and the contrastive baseline – this attention head function might be important in general. Next, in Section 4.1 we noticed clear difference between component contributions for first and later subwords – this suggests that some model components might be responsible for that.
In what follows, we are going to look at two attention head functions found earlier for machine translation, positional and subword merging heads, and see whether they are generally important. Differently from previous work, we define a head function based on token contributions within head and not attention weights as done previously. This is reasonable since attention weights might not reflect influences properly (Bastings & Filippova (2020); Kobayashi et al. (2020), among others).
Positional heads.
[…]
Figure 9a shows previous token heads in yellow. As in the earlier work for machine translation (Voita et al., 2019), we also see that for LLMs, (i) there are several previous tokens heads in the model, and (ii) almost all of them are by far the most important attention heads in the corresponding layers.
Subword merging heads.
[…]
We see subword merging heads in the bottom part of the network (Figure 9a, red), and are among the most important heads in the corresponding layers. This is interesting: previous work that noticed such head roles did not study their general relevance (Correia et al., 2019). Note that these subword merging heads are important for later subwords and not important otherwise – this explains the clusters we saw in Figure 8c.
We would like to highlight that for language models, we are the first to talk about general importance of attention heads and to find that some of the most important heads are previous token and subword merging heads. Future work might explore the functions of other important heads in the model.
Figure 10: Examples of important information flow subgraphs ($\tau=0.01$). Llama 2-7B.
4.4 Peculiar Information Flow Patterns, or Periods Acting as BOS
In Section 4.1 we talked about general contribution patterns and saw visible clusters corresponding to the input tokens’ part of speech. Now, let us go deeper and look in detail at one of the clusters. We choose the outlier punctuation cluster shown in yellow in Figure 8 (to the right) – this cluster corresponds to the first period in a text.777823 out of 826 datapoints.
Figure 9b shows the average importance of model components for examples in this cluster. We see that for these examples, the residual stream ignores all attention and FFN blocks in all the layers except for the first and last few: for most of the layers, contributions of all model components are near zero. When we look at the information flow graphs for these examples, we see that, even for a rather small threshold $\tau=0.01$, bottom-to-top processing happens largely via “untouched” residual connection (Figure 10). In Appendix C we show that up to the last three layers, this residual stream takes the role of the BOS token and future tokens treat this residual stream as such. Concurrent work (Cancedda, 2024) suggests that early FFN’s update into a particular subspace is responsible for this behavior. While for some examples periods acting as BOS might be reasonable, this also happens in cases where the period does not have an end-of-sentence role. For example, Figure 10 (right): while in a sentence For the second game in a row, St. Thomas ... the first period does not have the end-of-sentence meaning, the residual stream is still acting in the same manner. In future work, it might be valuable to explore whether this behavior might cause incorrect generation behavior.
5 Domain-Specific Model Components
Figure 11: Average importance of attention heads and FFNs for different datasets. For non-general domains, we show only heads with importance higher than 0.015. Llama 2-7B.
We just saw that some of attention head functions are universally important across all predictions. Now, let us see whether some model components are not important universally but are specialized for specific types of data.
Setting.
We run Llama 2-7B on 1000 sentences from different domains and evaluate the average importance of model components within each domain. Specifically, we consider (i) C4 (Raffel et al., 2020), (ii) different languages: English, Russian, Italian, Spanish FloRes-200 devtest sets (NLLB et al., 2022; Goyal et al., 2022), (iii) code data from CodeParrot [^8] https://huggingface.co/datasets/codeparrot/github-code, (iv) addition/subtraction data (Stolfo et al., 2023).
5.1 Components are Specialized
Figure 11 shows the results: generic importance of attention heads and FFN blocks as well as their importance for specific datasets.
Coarse grained: domains and languages.
From Figure 11 (left) we see that generally unimportant attention heads become highly relevant for specific tasks. For example, the most important heads for addition are among the lowest scoring heads when looking at C4 (Figure 11 (left), large yellow blobs to the left). Another interesting observation is that domain-specific heads are different across domains: important heads for addition, code, and non-English are not the same heads. Overall, we can conclude that model components are largely specialized.
Fine-grained: addition vs subtraction.
Figure 11 (right) shows more fine-grained results where instead of different domains, we look at tasks within the same narrow domain: addition vs subtraction. While the important heads for addition and subtraction largely intersect, we see several heads that are active only for one task and not the other (bright blue and yellow blobs). This means that this fine-grained specialization might be responsible for “reasoning” inside the model and not just domain-specific processing; future work may validate this further. Interestingly, we did not find similar specialization between languages and only find non-English heads – probably, portions of training data in these languages were not large enough to have dedicated language-specific heads.
Finally, we can also see the difference in the importance of the feed-forward blocks with respect to each domain. For example, the last FFN layer is much less relevant for non-English than other domains (Figure 11, left), the importance of FFN blocks for addition and subtraction falls to zero in some layers (Figure 11, right). In future work, one might conduct a more fine-grained analysis of the importance of feed-forward layers by looking at the relevance and functions of individual neurons (along the lines of Geva et al. (2021); Dai et al. (2022); Voita et al. (2023) but taking into account neuron importance for specific prediction).
Figure 12: Top 10 tokens after projecting singular values of $W_{OV}$ onto the unembedding matrix. Singular value ids are shown in gray.
5.2 Specialized Heads Output
Topic-Related Concepts
In this section, we analyze what our domain-specific heads write into the residual stream, and find that some of them write highly interpretable and topic-related concepts.
Weight matrices analysis with SVD.
As we illustrated in Figure 4, $W^{h}_{OV}$ transforms representations from each of the residual streams [^9] After applying layer normalization first. into vectors that are added to the current residual stream. To understand what kind of information is embedded in this transformation, we use Singular Value Decomposition (SVD).
[…]
${\bm{x}}W^{h}_{OV}=({\bm{x}}U\Sigma)V^{T}=\sum_{i=1}^{r}({\bm{x}}{\bm{u}}_{i}\sigma_{i}){\bm{v}}_{i}^{T}.$
Here, each ${\bm{u}}_{i}\sigma_{i}\in\mathbb{R}^{d\times 1}$ can be interpreted as a key that is compared to the query (${\bm{x}}$) via dot product (Molina, 2023). Each query-key dot-product weights the right singular vector ${\bm{v}}_{i}^{T}$. If we project these right singular vectors to the unembedding matrix (${\bm{v}}_{i}^{T}W_{U}$), we get an interpretation of the attention head’s influence in terms of concepts (i.e., tokens) it promotes in the residual stream.
Top singular values.
For some of the heads specific to code and non-English inputs we saw in Figure 11, Figure 12 shows the top 10 tokens that come from the described above projection. We see that code-specific heads promote tokens related to coding and technology in a more general sense: e.g., tokens related to Apple, like iOS, Xcode, iPhone, etc. Heads most active for non-English promote tokens in multiple languages, avoiding English ones. Also, we see tokens related to cities (Sydney), counties (Malay), and currencies (£). Overall, we looked at the functions of attention heads from two perspectives: (i) when a head is active, and (ii) how it updates the residual stream. While a similar kind of analysis was done before for neurons (Voita et al., 2023), our method made it possible to talk about entire model components being active for a prediction or not.
6 Related Work
Earlier works evaluating the importance of model components include Bau et al. (2019), who identified important neurons in NMT models shared across models, and Voita et al. (2019) who looked at the relevance of entire attention heads in the Transformer from two perspectives: attribution and pruning. More recent research focuses more on task-specific rather than overall importance. They aim to find subparts of LMs responsible for these tasks and rely on activation patching methodology, firstly introduced to analyze biases in LMs by Vig et al. (2020), and used across several others (Heimersheim & Janiak, 2023). Initial works discovered circuits for different tasks of GPT2, like IOI (Wang et al., 2023) and greater-than (Hanna et al., 2023). Patching has been also used to locate factual knowledge in LMs (Meng et al., 2022), and to discover task vectors behind in-context learning capabilities (Hendel et al., 2023; Todd et al., 2023). Our work can be related to the line of research developing methods to automatize the discovery of circuits implemented by LMs (Syed et al., 2023; Conmy et al., 2023).
7 Conclusions
We view computations inside the Transformer as information flowing between token representations through model components. Using this view, we propose to interpret language model predictions by extracting the important part of the overall information flow. Our method for extracting these important information flow routes is automatic, highly efficient, applicable to any prediction, more versatile and informative compared to existing pipelines.
8 Acknowledgments
We would like to thank Yihong Chen, Christoforos Nalmpantis, Igor Tufanov, Nicola Cancedda and Eduardo Sánchez for the insightful discussions.
Appendix A Linearizing the Layer Normalization
Given an input representation ${\bm{x}}_{j}$, the layernorm computes
$\text{LN}({\bm{x}}_{j})=\frac{{\bm{x}}_{j}-\mu({\bm{x}}_{j})}{\sigma({\bm{x}}_{j})}\odot\mathbf{\gamma}+\mathbf{\beta}$ (5)
with $\mu$ and $\sigma$ obtaining the mean and standard deviation, and $\gamma\in\mathbb{R}^{d}$ and $\beta\in\mathbb{R}^{d}$ refer to learned element-wise transformation and bias respectively. Considering $\sigma({\bm{x}}_{j})$ as a constant, LN can be treated as a constant affine transformation:
$\text{LN}({\bm{x}}_{j})={\bm{x}}_{j}L+\mathbf{\beta}$ (6)
where $L\in\mathbb{R}^{d\times d}$ represents a matrix that combines centering, normalizing, and scaling operations together.
${\mathbf{L}:=\frac{1}{\sigma({\bm{x}}_{j})}\left[\begin{array}[]{cccc}\gamma_{1}&0&\cdots&0\\ 0&\gamma_{2}&\cdots&0\\ \cdots&\cdots&\cdots&\cdots\\ 0&0&\cdots&\gamma_{n}\end{array}\right]\left[\begin{array}[]{cccc}\frac{n-1}{n}&-\frac{1}{n}&\cdots&-\frac{1}{n}\\ -\frac{1}{n}&\frac{n-1}{n}&\cdots&-\frac{1}{n}\\ \cdots&\cdots&\cdots&\cdots\\ -\frac{1}{n}&-\frac{1}{n}&\cdots&\frac{n-1}{n}\end{array}\right]}$
The linear map on the right subtracts the mean to the input vector, ${\bm{x}}_{j}^{\prime}={\bm{x}}_{j}-\mu({\bm{x}}_{j})$. The left matrix performs the hadamard product with the layer normalization weights (${\bm{x}}_{j}^{\prime}\odot\gamma$).
Appendix B Examples of Routes
Figures 13 and 14 show the information flow routes for IOI and greater-than tasks extracted for GPT2-small and OPT-125m.
Figure 13: The important information flow routes, IOI task, $\tau=0.04$.
Figure 14: The important information flow routes, greater-than task, $\tau=0.04$.
Appendix C Period Acting as BOS
Figure 15 shows an example of an attention map for one of the heads in Llama 2-7b. We see that after the first period, attention is spread between the BOS token and this period.
Figure 15: Llama 2 7B attention weights matrix L18H3.